Listed below are amounts (in millions of dollars) collected from parking meters by a security service company and other companies during similar time periods. Do the limited data listed here show evidence of 1.5 stealing by the security service companys employees? Security Service Company: 1.7 1.8 1.6 1.6 1.5 1.7 1.6 1.6 1.4 1.9 1.8 1.7 1.7 1.6 1.4 1.7 1.4 1.9 Other Companies Find the coefficient of variation for each of the two samples, then compare the variation. The coefficient of variation for the amount collected by the security service company is % (Round to one decimal place as needed.) The coefficient of variation for the amount collected by the other companies is % (Round to one decimal place as needed.) Do the limited data listed here show evidence of stealing by the security service companys employees? Consider a difference of greater than 1% to be significant. 0 A. No. There is a significant difference in the variation. O B. Yes. There is a significant difference in the variation. C. Yes. There is not a significant difference in the variation. O D. No. There is a not significant difference in the variation.
Answer #1
S S Company Other Company
1.4 1.9
1.7 1.8
1.6 1.6
1.5 1.7
1.6 1.6
1.4 1.9
1.8 1.7
1.7 1.6
1.4 1.7
1.5 1.8
x-bar = 1.56 1.73
sd = 0.142984071 0.115950181
CV = (sd/x-bar) * 100 = 9.165645551 6.702322594

(a) 9.2

(b) 6.7

(c) 6.7 * 1.1 = 7.37

Since 9.2 > 7.7, the answer is Option (B).